OpenStudy (anonymous):

A group of 100 people, some students and some faculty, attended a museum opening. Each student paid $10 per person for entrance to the museum and each of the faculty paid $25 per person for entrance. If the total paid, for all 100 people, was $1300, how many students attended the museum opening?

OpenStudy (texaschic101):

s = students f = faculty s + f = 100 ---> f = 100 - s 10s + 25f = 1300 can you solve it now ?

OpenStudy (texaschic101):

just sub in 100 - s in for f in the equation 10s + 25f = 1300 and you will be able to find the number of students (s)